# UlRe-NeRF: 3D Ultrasound Imaging through Neural Rendering with Ultrasound Reflection Direction Parameterization

### v1

@@ 1 Introduction
Figure 1: UlRe-NeRF for High-Fidelity 3D Ultrasound Imaging. This figure illustrates our proposed ultrasound neural rendering architecture, which integrates implicit neural representation and volume rendering to enhance 3D ultrasound imaging. The system comprises two main modules: the spatial MLP and the directional MLP. The spatial MLP predicts physical property parameters, such as attenuation coefficients, reflectivity, scattering density, and amplitude, through high-dimensional position encoding. This simulates the propagation and reflection behavior of ultrasound waves within the medium. The directional MLP processes the input ray directions, generating view-dependent intensity estimates that capture complex reflection phenomena.
@@
Ultra-NeRF is the first work to bring NeRF into the field of ultrasound[30]. It builds on the ultrasound volume rendering method by Salehi et al. and combines a coordinate-based MLP that maps from 3D coordinates to five physical parameters, achieving high-quality synthetic ultrasound images[19]. However, it is still view-independent, so cannot achieve ray interactions and the Fresnel Effect, and it does not allow rendering complex ultrasound artifacts, such as reverberations. To address these issues, we drew inspiration from Ref-NeRF and proposed a new model that considers both reflection intensity and volume rendering (Figures 1b and 1c). Specifically, we present a ultrasound neural rendering architecture based on implicit neural networks and volume rendering, optimizing the rendering process to better simulate the physical properties, artifacts, and deformations of the medium. We call this model UlRe-NeRF, which can accurately reproduce the propagation and reflection behavior of ultrasound waves within the medium, as depicted in Figure 1. In summary, our main contributions are as follows:
@@ 2.1 Variants of NeRF for quality
Neural Radiance Fields (NeRF) uses implicit representation and explicit rendering to achieve novel view synthesis[14]. Given any input 3D position $x$ and view direction $d$, NeRF uses a spatial MLP to output the density $\tau(x)$ and bottleneck vector $b(x)$ from 3D position $x$, and then use a second directional MLP to outputs the color of light $c(x,d)$ emitted by a particle at 3D position $x$ and direction $d$ (see Figure 4 for a visualization). The densities $\left\{\tau_{i}(x)\right\}$ and colors $\left\{c_{i}(x,d)\right\}$ queried along the camera ray $x_{i}=o+t_{i}d,d_{i}=d$ originating at $o$ with direction $d$, are alpha composited [11] to obtain the color of the pixel corresponding to the ray.
@@ 4.3 Ablation Study Results of the Rendering Components
5 Discussion & Conclusion
In this paper, we present UlRe-NeRF, a more effective and precise Neural Rendering method for 3D Ultrasound Imaging. We incorporate additional inductive biases about ultrasound: improving implicit representation network with reflection direction parameterization and Reflective Harmonic Encoding (RHE) to capture complex tissue reflections, improving explicit volumic rendering with integration of reflection intensity and volume rendering to simulate the medium’s physical properties and light interactions.
The experiments show that our approach can achieve high-fidelity 3D ultrasound imaging synthesizes: it learns view-dependent geometry and capture complex tissue structures; it simulates various physical property parameters of the the medium; it reproduces the propagation and reflection behavior of ultrasound waves in the medium.
Our study provides an important advance in ultrasound image reconstruction techniques, demonstrating the potential of ultrasound neural rendering architectures based on implicit neural networks and explicit ultrasound volume rendering to improve ultrasound image quality[15].
The characteristics of ultrasonic imaging itself limit the usage of hierarchical sampling strategy. Another methods specific for ultrasound to make rendering more sample-efficient is required. Although we have proposed new explicit ultrasound volume rendering methods to improve the contrast and clarity of physical properties simulation, further progress towards more efficient and reasonable methods is required to simulate the propagation and reflection behavior of ultrasound waves. Besides, for complex but more general freehand scanning situation, the ultrasound wave is much more sparse, in which additional regularization for view synthesis from sparse inputs is required.
We believe that this work will inspire further exploration of implicit representations in 3D Ultrasound for medical purpose.

### v2

@@ 1 Introduction
Figure 1: UlRe-NeRF for High-Fidelity 3D Ultrasound Imaging. This figure illustrates our proposed ultrasound neural rendering architecture, which integrates implicit neural representation (b) and volume rendering (c) to enhance 3D ultrasound imaging. The system comprises two main modules: the spatial MLP and the directional MLP. The spatial MLP predicts physical property parameters, such as attenuation coefficients, reflectivity, scattering density, and amplitude, through high-dimensional position encoding. This simulates the propagation and reflection behavior of ultrasound waves within the medium. The directional MLP processes the input ray directions, generating view-dependent intensity estimates that capture complex reflection phenomena.
@@
Ultra-NeRF is the first work to bring NeRF into the field of ultrasound[29]. It builds on the ultrasound volume rendering method by Salehi et al. and combines a coordinate-based MLP that maps from 3D coordinates to five physical parameters, achieving high-quality synthetic ultrasound images[18]. However, it is still view-independent, so cannot achieve ray interactions and the Fresnel Effect, and it does not allow rendering complex ultrasound artifacts, such as reverberations. To address these issues, we drew inspiration from Ref-NeRF and proposed a new model that considers both reflection intensity and volume rendering (Fig. 1(b) and 1(c)). Specifically, we present a ultrasound neural rendering architecture based on implicit neural networks and volume rendering, optimizing the rendering process to better simulate the physical properties, artifacts, and deformations of the medium. We call this model UlRe-NeRF, which can accurately reproduce the propagation and reflection behavior of ultrasound waves within the medium, as depicted in Fig. 1. In summary, our main contributions are as follows:
@@ 2.1 Variants of NeRF for quality
Neural Radiance Fields (NeRF) uses implicit representation and explicit rendering to achieve novel view synthesis[14]. Given any input 3D position $x$ and view direction $d$, NeRF uses a spatial MLP to output the density $\tau(x)$ and bottleneck vector $b(x)$ from 3D position $x$, and then use a second directional MLP to outputs the color of light $c(x,d)$ emitted by a particle at 3D position $x$ and direction $d$ (see Fig. 3 for a visualization). The densities $\left\{\tau_{i}(x)\right\}$ and colors $\left\{c_{i}(x,d)\right\}$ queried along the camera ray $x_{i}=o+t_{i}d,d_{i}=d$ originating at $o$ with direction $d$, are alpha composited [11] to obtain the color of the pixel corresponding to the ray.
@@ 4.3 Ablation Study Results of the Rendering Components
5 Discussion
In this paper, we present UlRe-NeRF, a more effective and precise Neural Rendering method for 3D Ultrasound Imaging. We incorporate additional inductive biases about ultrasound: enhancing the implicit representation network with reflection direction parameterization and Reflective Harmonic Encoding (RHE) to capture complex tissue reflections, and improving explicit volumetric rendering by integrating reflection intensity and volume rendering to simulate the medium’s physical properties and light interactions.
Our experiments show that UlRe-NeRF achieves high-fidelity 3D ultrasound imaging synthesis. It effectively learns view-dependent geometry, captures complex tissue structures, simulates various physical properties of the medium, and accurately reproduces the propagation and reflection behavior of ultrasound waves. This study marks an important advance in ultrasound image reconstruction techniques, demonstrating the potential of ultrasound neural rendering architectures based on implicit neural networks and explicit ultrasound volume rendering to improve image quality.
However, the inherent characteristics of ultrasonic imaging limit the usage of hierarchical sampling strategies. Thus, methods specifically designed for ultrasound rendering to enhance sample efficiency are required. Although we have introduced new explicit ultrasound volume rendering techniques to improve contrast and clarity in physical property simulations, further advancements are needed to develop more efficient and precise methods for simulating the propagation and reflection behavior of ultrasound waves. Additionally, in more complex freehand scanning situations, where ultrasound data is sparse, additional regularization for view synthesis from sparse inputs is necessary.
6 Conclusion
The UlRe-NeRF model represents a significant advancement in 3D ultrasound imaging by combining implicit neural networks with explicit volumetric rendering. Our approach successfully addresses the limitations of traditional ultrasound imaging methods, providing high-fidelity reconstructions of complex tissue structures and realistic simulations of wave propagation. While our work has made considerable progress, further research is needed to improve sample efficiency and address challenges associated with sparse data in freehand scanning scenarios. We anticipate that this study will inspire continued exploration and refinement of neural rendering techniques for medical applications, contributing to the future development of ultrasound imaging technology.